OpenStudy (raudhat24):

The total snowfall in Buffalo, NY one year was one hundred twenty inches. If the total snowfall the following year was ninety inches, what was the percent change? Show and explain your solution.

OpenStudy (anonymous):

25% change 120-90=30 30/120 = 1/4 = 25% i think

OpenStudy (raudhat24):

thanks